# RG-stable parameter relations of a scalar field theory in absence of a symmetry

### v3

@@ 2 A toy model with RG-stable parameter relations in the absence of a symmetry
$V_{R}=\tfrac{1}{2}m_{11}^{2}\,\left(\varphi_{1}^{2}\,-\,\varphi_{2}^{2}\right)+m_{12}^{2}\,\varphi_{1}\varphi_{2}+\tfrac{1}{24}\lambda_{1111}\,\left(\varphi_{1}^{4}+\varphi_{2}^{4}\right)+\tfrac{1}{4}\lambda_{1122}\,(\varphi_{1}\varphi_{2})^{2}+\tfrac{1}{6}\lambda_{1112}\,\left(\varphi_{1}^{2}\,-\,\varphi_{2}^{2}\right)\,\varphi_{1}\varphi_{2}\,,$
@@
$\mathscr{L}=\partial_{\mu}\Phi\partial^{\mu}\Phi^{*}-m_{1}^{2}\Phi^{*}\Phi-(m_{2}^{2}\Phi^{2}+{\rm c.c.})-\lambda_{1}(\Phi^{*}\Phi)^{2}-(\lambda_{2}\Phi^{4}+{\rm c.c.})-\bigl{(}\lambda_{3}\Phi^{2}+{\rm c.c.})\Phi^{*}\Phi\,,\phantom{xxx}$ (2.15)
@@
$\lambda_{2}$ $=$ $\tfrac{1}{96}\bigl{[}\lambda_{1111}+\lambda_{2222}-6\lambda_{1122}+3i(\lambda_{1112}-\lambda_{1222})\bigr{]}\,,$ (2.19)
$\lambda_{3}$ $=$ $\tfrac{1}{24}\bigl{[}\lambda_{1111}-\lambda_{2222}+2i(\lambda_{1112}+\lambda_{1222})\bigr{]}\,.$ (2.20)
@@ 3 Complexification of the toy model of two real scalar fields
[…]
@@
Similarly, $\Lambda_{{\bar{a}}{\bar{b}}{\bar{c}}{\bar{d}}}$ is symmetric under the permutation of the indices ${\bar{a}}{\bar{b}}{\bar{c}}{\bar{d}}$, $\Lambda_{a{\bar{b}}{\bar{c}}{\bar{d}}}$ is symmetric under the permutation of the indices ${\bar{b}}{\bar{c}}{\bar{d}}$, $\Lambda_{ab{\bar{c}}{\bar{d}}}$ is separately symmetric under the interchange of the indices $ab$ and ${\bar{c}}{\bar{d}}$, respectively [as indicated in eq. (3.6)], $\Lambda_{abc{\bar{d}}}$ is symmetric under the permutation of the indices $abc$, and $\Lambda_{abcd}$ is symmetric under the permutation of the indices $abcd$.
@@
$M_{1\bar{1}}^{2}=M_{2\bar{2}}^{2}\,,\qquad\qquad\,\,\,M_{1\bar{2}}=0\,,\qquad\qquad\quad\,\,M_{11}^{2}=-M_{22}^{2}\,,$ (3.13)
@@
where in eq. (3.16), we have made use of the last relation of eq. (3.8). This leaves us with five independent squared-mass parameters and 19 independent quartic coupling parameters. Observe that the last relation in eq. (3.13) and the two relations of eq. (3.14) [which exclusively depend on self coupling tensors with four unbarred indices] are analogous to the three relations given in eq. (2.2).
@@
$V_{C}$ $=$ $M^{2}\left(|\Phi_{1}|^{2}\,+|\Phi_{2}|^{2}\right)\,+\bigl{[}\bar{M}^{2}\left(\Phi_{1}^{2}-\Phi_{2}^{2}\right)+M^{2}_{12}\,\Phi_{1}\Phi_{2}+{\rm c.c.}\bigr{]}$ (3.17)
@@
$+\bigl{[}\Lambda_{10}(\Phi_{1}^{2}|\Phi_{2}|^{2}-\Phi^{2}_{2}|\Phi_{1}|^{2})+{\rm c.c.}\bigr{]}+\bigl{[}\Lambda_{11}\bigl{(}\Phi_{1}^{3}\Phi_{2}^{*}+\Phi_{2}^{3}\Phi_{1}^{*}\bigr{)}+{\rm c.c.}\bigr{]}\,.$
@@
so that all scalar potential coefficients are real, which finally leaves us with three independent real squared-mass parameters ($M^{2}$, $\bar{M}^{2}$, and $M_{12}^{2}$) and 11 independent real quartic coupling parameters ($\Lambda_{i}$ for $i=1,2,\ldots,11$) that govern the complexification of the toy model of Section 2. In particular, note that
@@ 4 One-loop beta functions of the complexified model
$\beta^{I}_{\Lambda_{{\bar{a}}{\bar{b}}{\bar{c}}{\bar{d}}}}$ $=$ $\frac{1}{96}\sum_{\rm perm}\Lambda_{{\bar{a}}{\bar{b}}{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{c}}{\bar{d}}}+\frac{1}{256}\sum_{\rm perm}\Lambda_{e{\bar{a}}{\bar{b}}{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{c}}{\bar{d}}}$ (4.18)
$=$ $\frac{1}{12}(\Lambda_{{\bar{a}}{\bar{b}}{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{c}}{\bar{d}}}+\Lambda_{{\bar{a}}{\bar{c}}{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{b}}{\bar{d}}}+\Lambda_{{\bar{a}}{\bar{d}}{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{b}}{\bar{c}}})+\frac{1}{32}(\Lambda_{e{\bar{a}}{\bar{b}}{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{c}}{\bar{d}}}+\Lambda_{e{\bar{a}}{\bar{c}}{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{b}}{\bar{d}}}+\Lambda_{e{\bar{a}}{\bar{d}}{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{b}}{\bar{c}}}),\phantom{xxxx}$
$\beta^{I}_{\Lambda_{ab{\bar{c}}{\bar{d}}}}$ $=$ $\frac{3}{2}\Lambda_{abef}\Lambda_{{\bar{e}}{\bar{f}}{\bar{c}}{\bar{d}}}+\frac{1}{24}\bigl{(}\Lambda_{ab{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{c}}{\bar{d}}}+\Lambda_{ae{\bar{f}}{\bar{c}}}\Lambda_{bf{\bar{e}}{\bar{d}}}+\Lambda_{be{\bar{f}}{\bar{c}}}\Lambda_{af{\bar{e}}{\bar{d}}}+\Lambda_{ae{\bar{f}}{\bar{d}}}\Lambda_{bf{\bar{e}}{\bar{c}}}+\Lambda_{be{\bar{f}}{\bar{d}}}\Lambda_{af{\bar{e}}{\bar{c}}}\bigr{)}$ (4.19)
$\qquad+\frac{3}{32}\bigl{(}2\Lambda_{abe{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{c}}{\bar{d}}}+\Lambda_{aef{\bar{c}}}\Lambda_{b{\bar{e}}{\bar{f}}{\bar{d}}}+\Lambda_{bef{\bar{c}}}\Lambda_{a{\bar{e}}{\bar{f}}{\bar{d}}}+\Lambda_{aef{\bar{d}}}\Lambda_{b{\bar{e}}{\bar{f}}{\bar{c}}}+\Lambda_{bef{\bar{d}}}\Lambda_{a{\bar{e}}{\bar{f}}{\bar{c}}}\bigr{)}\,,$
$\beta^{I}_{\Lambda_{abc{\bar{d}}}}$ $=$ $\frac{1}{24}\bigl{(}\Lambda_{ab{\bar{e}}{\bar{f}}}\Lambda_{efc{\bar{d}}}+\Lambda_{ac{\bar{e}}{\bar{f}}}\Lambda_{efb{\bar{d}}}+\Lambda_{bc{\bar{e}}{\bar{f}}}\Lambda_{efa{\bar{d}}}\bigr{)}+\frac{1}{12}\bigl{(}\Lambda_{abe{\bar{f}}}\Lambda_{cf{\bar{e}}{\bar{d}}}+\Lambda_{ace{\bar{f}}}\Lambda_{bf{\bar{e}}{\bar{d}}}+\Lambda_{bce{\bar{f}}}\Lambda_{af{\bar{e}}{\bar{d}}}\bigr{)}$ (4.20)
$\qquad+\frac{1}{4}\bigl{(}\Lambda_{abef}\Lambda_{c{\bar{e}}{\bar{f}}{\bar{d}}}+\Lambda_{acef}\Lambda_{b{\bar{e}}{\bar{f}}{\bar{d}}}+\Lambda_{bcef}\Lambda_{a{\bar{e}}{\bar{f}}{\bar{d}}}\bigr{)}\,.$
@@
Following our one-loop analysis, we shall consider the two-loop beta function for the complexified model. If we follow our previous technique, we should rewrite eq. (2.6) in terms of the parameters $M^{2}_{{\bar{c}}{\bar{d}}}$, $M_{{\bar{c}}d}$, $\Lambda_{cd{\bar{a}}{\bar{b}}}$, $\Lambda_{{\bar{a}}{\bar{b}}{\bar{c}}{\bar{d}}}$, and $\Lambda_{d{\bar{a}}{\bar{b}}{\bar{e}}}$. However, we can identify the possible index structure of the various terms. Similar to eqs. (5.2)–(5.4), one can derive separate relations that must be satisfied if the beta function vanishes. Since the complexified theory does possess a symmetry that guarantees the relations exhibited in eqs. (3.13)–(3.16), we are assured that the two-loop beta function will vanish. Among all the relations obtained, we find that
@@ 6 Complexification and realification revisited
[…]
@@ 7 Summary of Results
where the $f_{k}(M^{2},\Lambda)$ are functions of the squared-mass and self coupling parameters. Each term in the sum will contain one factor of $M^{2}$ and $n$ factors of $\Lambda$ at order $n$ in the perturbation expansion. The tensor $M^{2}$ can have index structure $cd$, ${\bar{c}}{\bar{d}}$, or $c{\bar{d}}$, and the tensor $\Lambda$ can have index structure $cdef$, $cde{\bar{f}}$, $cd{\bar{e}}{\bar{f}}$, $c{\bar{d}}{\bar{e}}{\bar{f}}$, or ${\bar{c}}{\bar{d}}{\bar{e}}{\bar{f}}$. By appropriate choices of the index structure along with some appropriate Kronecker deltas to tie together some unbarred/barred index pairs, the index structure of the $f_{k}$ must be ${\bar{a}}{\bar{b}}$ as indicated in eq. (7.2). As a simple example, at one-loop order eq. (7.2) takes the form
@@ Appendix A Beta functions of a toy model of two real scalar fields
$\beta^{I}_{m_{22}^{2}}$ $=$ $m_{22}^{2}\lambda_{2222}+m_{11}^{2}\lambda_{1122}^{2}+2m_{12}^{2}\lambda_{1222}\,.$ (A.3)
@@
where “perm” indicates that the sum includes terms in which the uncontracted indices $i$, $j$, $k$, and $\ell$ have been permuted in all possible ways. In addition, there are implicit sums over each repeated index pair. Then, we obtain
@@
Next, we record the following results:
$C_{1111}=\lambda_{1111}^{3}+\lambda_{1111}\bigl{(}4\lambda_{1112}^{2}+3\lambda_{1122}^{2}+\lambda_{1222}^{2}\bigr{)}+3\lambda_{1112}\lambda_{1122}(\lambda_{1112}+\lambda_{1222})+\lambda_{1112}\lambda_{1222}\lambda_{2222}\,,$ (A.22)
$C_{2222}=\lambda_{2222}^{3}+\lambda_{2222}\bigl{(}4\lambda_{1222}^{2}+3\lambda_{1122}^{2}+\lambda_{1112}^{2}\bigr{)}+3\lambda_{1222}\lambda_{1122}(\lambda_{1112}+\lambda_{1222})+\lambda_{1111}\lambda_{1112}\lambda_{1222}\,,$ (A.23)
$D_{1111}=\lambda_{1111}^{3}+\lambda_{1111}\bigl{(}4\lambda_{1112}^{2}+\lambda_{1122}^{2}\bigr{)}+2\lambda_{1122}^{3}+\lambda_{1122}\bigl{(}\lambda_{1222}^{2}+2\lambda_{1112}\lambda_{1222}+5\lambda_{1112}^{2}\bigr{)}\,,$ (A.24)
$D_{2222}=\lambda_{2222}^{3}+\lambda_{2222}\bigl{(}4\lambda_{1222}^{2}+\lambda_{1122}^{2}\bigr{)}+2\lambda_{1122}^{3}+\lambda_{1122}\bigl{(}\lambda_{1112}^{2}+2\lambda_{1112}\lambda_{1222}+5\lambda_{1222}^{2}\bigr{)}\,.$ (A.25)

### v6

@@ 2 A toy model with RG-stable parameter relations in the absence of a symmetry
$\hskip-7.22743ptV_{R}=\tfrac{1}{2}m_{11}^{2}\,\left(\varphi_{1}^{2}\,-\,\varphi_{2}^{2}\right)+m_{12}^{2}\,\varphi_{1}\varphi_{2}+\tfrac{1}{24}\lambda_{1111}\,\left(\varphi_{1}^{4}+\varphi_{2}^{4}\right)+\tfrac{1}{4}\lambda_{1122}\,(\varphi_{1}\varphi_{2})^{2}+\tfrac{1}{6}\lambda_{1112}\,\left(\varphi_{1}^{2}\,-\,\varphi_{2}^{2}\right)\,\varphi_{1}\varphi_{2}\,$,
@@
[…]
@@
$\lambda_{2}$ $=$ $\tfrac{1}{96}\bigl[\lambda_{1111}+\lambda_{2222}-6\lambda_{1122}-4i(\lambda_{1112}-\lambda_{1222})\bigr]\,$, (2.19)
$\lambda_{3}$ $=$ $\tfrac{1}{24}\bigl[\lambda_{1111}-\lambda_{2222}-2i(\lambda_{1112}+\lambda_{1222})\bigr]\,.$ (2.20)
@@ 3 Complexification of the toy model of two real scalar fields
$\Phi_{1}\to-\Phi_{1}\quad\mbox{and}\quad\Phi_{2}\to-\Phi_{2}\quad\mbox{(taken simultaneously)},$ (3.4)
@@
Similarly, $\Lambda_{{\bar{a}}{\bar{b}}{\bar{c}}{\bar{d}}}$ is symmetric under the permutation of the indices ${\bar{a}}{\bar{b}}{\bar{c}}{\bar{d}}$, $\Lambda_{a{\bar{b}}{\bar{c}}{\bar{d}}}$ is symmetric under the permutation of the indices ${\bar{b}}{\bar{c}}{\bar{d}}$, $\Lambda_{ab{\bar{c}}{\bar{d}}}$ is separately symmetric under the interchange of the indices $ab$ and ${\bar{c}}{\bar{d}}$, respectively [as indicated in eq. (3.6)], $\Lambda_{abc{\bar{d}}}$ is symmetric under the permutation of the indices abcabc, and $\Lambda_{abcd}$ is symmetric under the permutation of the indices abcdabcd.
@@
$M_{1\bar{1}}^{2}=M_{2\bar{2}}^{2}\,,\qquad\qquad\operatorname{Re}M^{2}_{1\bar{2}}=0\,,\qquad\qquad\,\,\,M_{11}^{2}=-M_{22}^{2}\,$, (3.13)
@@
where in eq. (3.16), we have made use of the last relation of eq. (3.8). This leaves us with six independent squared-mass parameters and 19 independent quartic coupling parameters. Observe that the last relation in eq. (3.13) and the two relations of eq. (3.14) [which exclusively depend on self coupling tensors with four unbarred indices] are analogous to the three relations given in eq. (2.2).
@@
$V_{C}$ $=$ $M^{2}\left(|\Phi_{1}|^{2}\,+|\Phi_{2}|^{2}\right)+i\operatorname{Im}M^{2}_{1\bar{2}}\bigl(\Phi_{1}^{*}\Phi_{2}-\Phi_{1}\Phi_{2}^{*}\bigr)+\bigl[\bar{M}^{2}\left(\Phi_{1}^{2}-\Phi_{2}^{2}\right)+M^{2}_{12}\,\Phi_{1}\Phi_{2}+{\rm c.c.}\bigr]$ (3.17)
@@
$+\bigl[\Lambda_{10}(\Phi_{1}^{2}|\Phi_{2}|^{2}-\Phi^{2}_{2}|\Phi_{1}|^{2})+{\rm c.c.}\bigr]+\bigl[\Lambda_{11}\bigl(\Phi_{1}^{3}\Phi_{2}^{*}+\Phi_{2}^{3}\Phi_{1}^{*}\bigr)+{\rm c.c.}\bigr]\,$,
where $M^{2}\equiv M^{2}_{1\bar{1}}=M^{2}_{2\bar{2}}$ and $\bar{M}^{2}\equiv M_{11}^{2}=-M^{2}_{22}$.
@@
so that all scalar potential coefficients are real. It then follows that $M^{2}_{1\bar{2}}=0$, which finally leaves us with three independent real squared-mass parameters ($M^{2}$, $\bar{M}^{2}$, and $M_{12}^{2}$) and 11 independent real quartic coupling parameters ($\Lambda_{i}$ for $i=1,2,\ldots,11$) that govern the complexification of the toy model of Section 2. In particular, note that
@@ 4 One-loop beta functions of the complexified model
$\beta^{I}_{\Lambda_{{\bar{a}}{\bar{b}}{\bar{c}}{\bar{d}}}}$ $=$ $\sum_{\rm perm}\Lambda_{{\bar{a}}{\bar{b}}{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{c}}{\bar{d}}}+\frac{3}{8}\sum_{\rm perm}\Lambda_{e{\bar{a}}{\bar{b}}{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{c}}{\bar{d}}}$ (4.18)
$=$ $8(\Lambda_{{\bar{a}}{\bar{b}}{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{c}}{\bar{d}}}+\Lambda_{{\bar{a}}{\bar{c}}{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{b}}{\bar{d}}}+\Lambda_{{\bar{a}}{\bar{d}}{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{b}}{\bar{c}}})+3(\Lambda_{e{\bar{a}}{\bar{b}}{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{c}}{\bar{d}}}+\Lambda_{e{\bar{a}}{\bar{c}}{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{b}}{\bar{d}}}+\Lambda_{e{\bar{a}}{\bar{d}}{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{b}}{\bar{c}}}),\phantom{xxxx}$
$\beta^{I}_{\Lambda_{ab{\bar{c}}{\bar{d}}}}$ $=$ $144\Lambda_{abef}\Lambda_{{\bar{e}}{\bar{f}}{\bar{c}}{\bar{d}}}+4\bigl(\Lambda_{ab{\bar{e}}{\bar{f}}}\Lambda_{ef{\bar{c}}{\bar{d}}}+\Lambda_{ae{\bar{f}}{\bar{c}}}\Lambda_{bf{\bar{e}}{\bar{d}}}+\Lambda_{be{\bar{f}}{\bar{c}}}\Lambda_{af{\bar{e}}{\bar{d}}}+\Lambda_{ae{\bar{f}}{\bar{d}}}\Lambda_{bf{\bar{e}}{\bar{c}}}+\Lambda_{be{\bar{f}}{\bar{d}}}\Lambda_{af{\bar{e}}{\bar{c}}}\bigr)$ (4.19)
$\qquad+9\bigl(2\Lambda_{abe{\bar{f}}}\Lambda_{f{\bar{e}}{\bar{c}}{\bar{d}}}+\Lambda_{aef{\bar{c}}}\Lambda_{b{\bar{e}}{\bar{f}}{\bar{d}}}+\Lambda_{bef{\bar{c}}}\Lambda_{a{\bar{e}}{\bar{f}}{\bar{d}}}+\Lambda_{aef{\bar{d}}}\Lambda_{b{\bar{e}}{\bar{f}}{\bar{c}}}+\Lambda_{bef{\bar{d}}}\Lambda_{a{\bar{e}}{\bar{f}}{\bar{c}}}\bigr)\,$,
$\beta^{I}_{\Lambda_{abc{\bar{d}}}}$ $=$ $4\bigl(\Lambda_{ab{\bar{e}}{\bar{f}}}\Lambda_{efc{\bar{d}}}+\Lambda_{ac{\bar{e}}{\bar{f}}}\Lambda_{efb{\bar{d}}}+\Lambda_{bc{\bar{e}}{\bar{f}}}\Lambda_{efa{\bar{d}}}\bigr)+8\bigl(\Lambda_{abe{\bar{f}}}\Lambda_{cf{\bar{e}}{\bar{d}}}+\Lambda_{ace{\bar{f}}}\Lambda_{bf{\bar{e}}{\bar{d}}}+\Lambda_{bce{\bar{f}}}\Lambda_{af{\bar{e}}{\bar{d}}}\bigr)$ (4.20)
$\qquad+24\bigl(\Lambda_{abef}\Lambda_{c{\bar{e}}{\bar{f}}{\bar{d}}}+\Lambda_{acef}\Lambda_{b{\bar{e}}{\bar{f}}{\bar{d}}}+\Lambda_{bcef}\Lambda_{a{\bar{e}}{\bar{f}}{\bar{d}}}\bigr)\,.$
@@
The structure of the expressions for the beta functions obtained in eqs. (4.6), (4.7), and (4.18)–(4.20) is noteworthy. In particular, covariance with respect to the unbarred/barred indices is respected by the beta function equations. That is, after carrying out the implicit summation over unbarred/barred index pairs, the index structure the beta function must match the corresponding free indices appearing on the right-hand side of the beta function expression. For example, the free indices on both sides of eq. (4.18) consist of four barred indices. The covariance of the other beta function equations can be easily verified.
@@
Following our one-loop analysis, we shall consider the two-loop beta function for the complexified model. If we follow our previous technique, we should rewrite eq. (2.6) in terms of the parameters $M^{2}_{{\bar{c}}{\bar{d}}}$, $M_{{\bar{c}}d}$, $\Lambda_{cd{\bar{a}}{\bar{b}}}$, $\Lambda_{{\bar{a}}{\bar{b}}{\bar{c}}{\bar{d}}}$, and $\Lambda_{d{\bar{a}}{\bar{b}}{\bar{e}}}$. However, we can identify the possible index structures of the various terms that respect the covariance properties of the corresponding beta functions. Similar to eqs. (5.2)–(5.4), one can derive separate relations that must be satisfied if the beta function vanishes. Since the complexified theory does possess a symmetry that guarantees the relations exhibited in eqs. (3.13)–(3.16), we are assured that the two-loop beta function will vanish. Among all the relations obtained, we find that
@@ 6 Complexification and realification revisited
The above procedure suggests an algorithm for constructing examples of real scalar field theories with an RG-stable parameter relation without a symmetry to explain its RG-stability. Start with a theory of $n$ complex scalars with parameter relations whose RG-stability can be accounted for by the symmetries of the model. From this theory, construct the corresponding theory of $n$ real scalars whose complexification yields the theory of $n$ complex scalars (using the method outlined above). If the symmetries of the complexified theory involve some symmetry transformation group that cannot be embedded in O($n)$, then such symmetries cannot survive as a legitimate symmetry of the theory of $n$ real scalars. This is precisely what happened in Section 3, where the symmetry employed [eq. (3.12)] corresponded to the matrix $\left(\begin{smallmatrix}\phantom{-}0&i\\ -i&0\end{smallmatrix}\right)$, which is not an element of O(2). Nevertheless, the vanishing of the beta functions of the parameter relations of the theory of $n$ complex scalars will still ensure the vanishing of the corresponding beta functions of the theory of $n$ real scalars, as discussed in Section 5.
@@ 7 Summary of Results
where the $f_{k}(M^{2},\Lambda)$ are functions of the squared-mass and self coupling parameters. Each term in the sum will contain one factor of $M^{2}$ and $n$ factors of $\Lambda$ at order $n$ in the perturbation expansion. The tensor $M^{2}$ can have index structure $cd$, ${\bar{c}}{\bar{d}}$, or $c{\bar{d}}$, and the tensor $\Lambda$ can have index structure cdefcdef, $cde{\bar{f}}$, $cd{\bar{e}}{\bar{f}}$, $c{\bar{d}}{\bar{e}}{\bar{f}}$, or ${\bar{c}}{\bar{d}}{\bar{e}}{\bar{f}}$. By appropriate choices of the index structure along with some appropriate Kronecker deltas to tie together some unbarred/barred index pairs, the index structure of the $f_{k}$ must be ${\bar{a}}{\bar{b}}$ as indicated in eq. (7.2). As a simple example, at one-loop order eq. (7.2) takes the form
@@ Acknowledgments
Additional acknowledgments
We thank Patrick Fox for pointing out the error in the original published version of eq. (3.13) and for other insightful remarks in discussions with H.E.H. at the Munich Institute for Astro-, Particle and BioPhysics (MIAPbP), which is funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – EXC-2094–390783311, and at the Aspen Center for Physics, which is supported by the National Science Foundation Grant No. PHY-2210452. H.E.H also greatly appreciates the hospitality and housing support of the NYU Center for Cosmology and Particle Physics during two recent visits, where he benefited from stimulating conversations with Ken Van Tilburg and his collaborators concerning the use of the spurion technique in determining RG fixed points.
@@ Appendix A Beta functions of a toy model of two real scalar fields
$\beta^{I}_{m_{22}^{2}}$ $=$ $m_{22}^{2}\lambda_{2222}+m_{11}^{2}\lambda_{1122}+2m_{12}^{2}\lambda_{1222}\,.$ (A.3)
@@
where “perm” indicates that the sum includes terms in which the uncontracted indices $i$, $j$, $k$, and $\ell$ have been permuted in all possible ways. In addition, there are implicit sums over each repeated index pair. For example, in light of eq. (2.6),
$\beta^{II}_{m_{11}^{2}+m_{22}^{2}}=\tfrac{1}{6}\bigl(A_{11}+A_{22}\bigr)-2\bigl(B_{11}+B_{22}\bigr)\,.$ (A.17)
We then obtain
@@
It is instructive to use the above results to compute the beta function of the parameter $m_{1}^{2}$ that appears in eq. (2.15). In light of eq. (2.16), it follows that $\beta_{m_{1}^{2}}=\tfrac{1}{2}\beta_{m_{11}^{2}+m_{22}^{2}}$. We now invert the results of eqs. (2.16)–(2.20) to obtain
$m_{11}^{2}=m_{1}^{2}+2\operatorname{Re}m_{2}^{2}\,,\qquad\quad m_{22}^{2}=m_{1}^{2}-2\operatorname{Re}m_{2}^{2}\,,\qquad\quad m^{2}_{12}=-2\operatorname{Im}m_{2}^{2}\,$, (A.23)
and
$\lambda_{1111}$ $=$ $6\lambda_{1}+12\operatorname{Re}(\lambda_{2}+\lambda_{3})\,$, (A.24)
$\lambda_{1112}$ $=$ $-12\operatorname{Im}\lambda_{2}-6\operatorname{Im}\lambda_{3}\,$, (A.25)
$\lambda_{1122}$ $=$ $2\lambda_{1}-12\lambda_{2}\,$, (A.26)
$\lambda_{1222}$ $=$ $12\operatorname{Im}\lambda_{2}-6\operatorname{Im}\lambda_{3}\,$, (A.27)
$\lambda_{2222}$ $=$ $6\lambda_{1}+12\operatorname{Re}(\lambda_{2}-\lambda_{3})\,.$ (A.28)
Using eqs. (A.4) and (A.17)–(A.21), one can obtain the one-loop and two-loop contributions to $\beta_{m_{1}^{2}}$,
$\beta^{I}_{m_{1}^{2}}$ $=$ $8\lambda_{1}m_{1}^{2}+24\operatorname{Re}(\lambda_{3}^{*}m_{2}^{2})\,$, (A.29)
$\beta^{II}_{m_{1}^{2}}$ $=$ $-8m_{1}^{2}\bigl(11\lambda_{1}^{2}+132|\lambda_{2}|^{2}+33|\lambda_{3}|^{2}\bigr)-528\lambda_{1}\operatorname{Re}(\lambda_{3}^{*}m_{2}^{2})-1056\operatorname{Re}(\lambda_{2}^{*}\lambda_{3}m_{2}^{2})\,$, (A.30)
after making use of eq. (2.16). Note that eq. (A.29) can also be derived directly from eq. (4.7). As expected, $\beta^{I}_{m_{1}^{2}}=\beta^{II}_{m_{1}^{2}}=0$ after imposing the relations $m_{1}^{2}=\lambda_{3}=0$ as indicated in eq. (2.22).
The structure of the expressions on the right-hand sides of eqs. (A.29) and (A.30) is noteworthy. Using the unbarred/barred index notation introduced in Section 3, one can write
$m_{1}^{2}$ $\equiv$ $M^{2}_{1\bar{1}}\,,\qquad m_{2}^{2}\equiv M^{2}_{\bar{1}\bar{1}}\,,\qquad(m_{2}^{2})^{*}\equiv M^{2}_{11}\,$, (A.31)
$\lambda_{1}$ $\equiv$ $\Lambda_{11\bar{1}\bar{1}}\,,\qquad\lambda_{2}\equiv\Lambda_{\bar{1}\bar{1}\bar{1}\bar{1}}\,,\qquad\,\,\,\lambda_{2}^{*}\equiv\Lambda_{1111}\,,\qquad\lambda_{3}\equiv\Lambda_{1\bar{1}\bar{1}\bar{1}}\,,\qquad\lambda_{3}^{*}\equiv\Lambda_{111\bar{1}}\,.$ (A.32)
In light of the covariance properties of the corresponding beta functions noted at the end of Section 4, it follows that the index structure of $\beta_{m_{1}^{2}}$ must contain an equal number of unbarred and barred indices. One can check that the expressions on the right-hand sides of eqs. (A.29) and (A.30) satisfy this requirement. Indeed, this requirement is equivalent to the result of the spurion analysis of Ref. [49]. [^11] We thank Ken Van Tilburg for communicating the results of Ref. [49] to us prior to its release. Such an analysis fixes the forms that can appear in the computation of $\beta_{m_{1}^{2}}$ at any loop order, and implies that eq. (2.22) is satisfied to all orders in perturbation theory.
Next, we note that in light of eq. (2.7),
$\beta^{II}_{\lambda_{1111}-\lambda_{2222}}$ $=$ $\tfrac{1}{3}\bigl(C_{1111}-C_{2222}\bigr)-6\bigl(D_{1111}-D_{2222})\,$, (A.33)
$\beta^{II}_{\lambda_{1112}+\lambda_{1222}}$ $=$ $\tfrac{1}{3}\bigl(C_{1112}+C_{1222}\bigr)-6\bigl(D_{1112}+D_{1222})\,.$ (A.34)
Thus, we record the following results:
$\hskip-36.135ptC_{1111}=\lambda_{1111}^{3}+\lambda_{1111}\bigl(4\lambda_{1112}^{2}+3\lambda_{1122}^{2}+\lambda_{1222}^{2}\bigr)+3\lambda_{1112}\lambda_{1122}(\lambda_{1112}+\lambda_{1222})+\lambda_{1112}\lambda_{1222}\lambda_{2222}\,$, (A.35)
$\hskip-36.135ptC_{2222}=\lambda_{2222}^{3}+\lambda_{2222}\bigl(4\lambda_{1222}^{2}+3\lambda_{1122}^{2}+\lambda_{1112}^{2}\bigr)+3\lambda_{1222}\lambda_{1122}(\lambda_{1112}+\lambda_{1222})+\lambda_{1111}\lambda_{1112}\lambda_{1222}\,$, (A.36)
$\hskip-36.135ptD_{1111}=\lambda_{1111}^{3}+\lambda_{1111}\bigl(4\lambda_{1112}^{2}+\lambda_{1122}^{2}\bigr)+2\lambda_{1122}^{3}+\lambda_{1122}\bigl(\lambda_{1222}^{2}+2\lambda_{1112}\lambda_{1222}+5\lambda_{1112}^{2}\bigr)\,$, (A.37)
$\hskip-36.135ptD_{2222}=\lambda_{2222}^{3}+\lambda_{2222}\bigl(4\lambda_{1222}^{2}+\lambda_{1122}^{2}\bigr)+2\lambda_{1122}^{3}+\lambda_{1122}\bigl(\lambda_{1112}^{2}+2\lambda_{1112}\lambda_{1222}+5\lambda_{1222}^{2}\bigr)\,.$ (A.38)